OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of 1/3 and passes through the point (-6, 0).

jimthompson5910 (jim_thompson5910):

point slope form is y - y1 = m(x - x1) where m is the slope (x1,y1) is the point the line goes through

jimthompson5910 (jim_thompson5910):

in this case m = 1/3 x1 = -6 y1 = 0